[ members of the computer club were assessed equal amounts to raise $1200 to buy new software. When 8 [NEW] members joined, the per-member assessment was reduced by $7.50. What [IS] the new size of the club? ]
Answer by TimothyLamb(4379)

1200 = (p - 7.50)(n + 8)
1200 = pn
p = 1200/n
---
1200 = (p - 7.50)(n + 8)
1200 = pn + 8p - 7.50n - 60
1260 = pn + 8p - 7.50n
1260 = (1200/n)n + 8(1200/n) - 7.50n
1260 = 1200 + 9600/n - 7.50n
60 = 9600/n - 7.50n
60n = 9600 - 7.50nn
7.50n^2 + 60n - 9600 = 0
---
the above quadratic equation is in standard form, with a=7.5, b=60, and c=-9600
---
to solve the quadratic equation, plug this:
7.5 60 -9600
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
n = 32
n = -40
---
the negative root doesn't make sense for the size of the club (number of members), so use the positive root
---
Answer:
the old size of the club is 32 members
the new size of the club is 40 members
---
the old per-member assessment was: 1200/32 = $37.50
the new per-member assessment is: 1200/40 = $30.00
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php